OpenStudy (naga447):

I need help with this equation 4 - (x+2) < -3(x+4) I keep getting the same awnser and I know that it is worng here's my work... 4-(x+2) < -3(x+4) first I use the distributive property 4x+2 < -3x+4 then I combine like terms 4x < -3x+6 then combine like terms again x < 6 but this is a multiple choice question and what I get is not on there. Could some one please explain what I am doing wrong?

OpenStudy (anonymous):

mistake in the first line

OpenStudy (anonymous):

you wrote "distribute" but you did not actually distribute, you rewrote with no parentheses

OpenStudy (naga447):

okay that makes sense! could you please enplane how I would do it correctly?

OpenStudy (anonymous):

\[4-(x+2) < -3(x+4)\] is the first line right?

OpenStudy (naga447):

Yeah

OpenStudy (anonymous):

when you distribute on the right hand sides you get \[-3(x+4)=-3x-12\]

OpenStudy (anonymous):

on the left hand side the stuff in the parentheses is not multiplied by 4 you need only distribute the minus sign \[4-(x+2)=4-x-2=2-x\]

OpenStudy (anonymous):

so after distributing and combining like terms you should be looking at \[2-x<-3x-12\]

OpenStudy (naga447):

okay so then I add -2 to both sides right?

OpenStudy (anonymous):

i would attack the variables first since you have \(-x\) on the left and \(-3x\) on the right i would add \(3x\) to both sides

OpenStudy (naga447):

2-4x < -12?

OpenStudy (naga447):

or would I add 3x to its self and get 2-4x < 6x - 12

OpenStudy (naga447):

@satellite73

OpenStudy (naga447):

oh its neither of thous...

OpenStudy (naga447):

I've got it its -10!!! x = -10